OpenStudy (anonymous):

Our parabola runs through (3,5) and also through (5,5), and it has of 2. What is the vertex form(s) of this quadratic equation? Draw a sketch if necessary to help you out.( |a|- is absolute value of a)

OpenStudy (anonymous):

it has (BLANK) of 2..... what is (BLANK)?

OpenStudy (anonymous):

the a value, sorry

OpenStudy (anonymous):

Well our equation is y=2(x-h) + k Since a = 2, right?

OpenStudy (anonymous):

yes

OpenStudy (ranga):

You can start with y = ax^2 + bx + c where a = 2 You are given two points. So you can solve for b and c. Then you can complete the square and put it in the vertex form.

OpenStudy (ranga):

y = 2x^2 + bx +c ---- (1) (3,5) and (5,5) are points on the parabola. when x = 3, y = 5. Substitute in (1) 5 = 2(3)^2 + 3b + c 5 = 18 + 3b + c 3b + c = -13 ----- (2) when x = 5, y = 5. Substitute in (1) 5 = 2(5)^2 + 5b + c 5 = 50 + 5b + c 5b + c = -45 ------ (3) Subtract (2) from (3) 2b = -45 + 13 = -32 b = -16 Put it in (2): 3(-16) + c = -13 c = -13 + 48 = 35 The parabola is: y = 2x^2 -16x + 35 Complete the square to put the equation in vertex form: y = 2x^2 -16x + 35 = 2(x^2 - 8x + 35/2) y = 2{ (x-4)^2 - 16 + 35/2 } y = 2 { (x-4)^2 + (-32 + 35)/2 } y = 2 { (x-4)^2 + (3)/2 } y = 2(x-4)^2 + 3 Compare this to the standard equation of a parabola in vertex form: a(x-h)^2 + k where (h,k) is the vertex and you can find the vertex of the parabola y = 2(x-4)^2 + 3

OpenStudy (ranga):

But vertex is not asked. Just the vertex form of the parabola.